In an electric lamp, 2.50C pass through each second. How many electrons does this represent?


1
Expert's answer
2022-03-20T18:48:40-0400

The number of electrons passing through the lamp

"N=\frac{q}{e}=\frac{2.5}{1.6\times 10^{-19}}=1.56\times 10^{19}"